Many random number generators allow users to specify the range of the random numbers to be produced. Suppose that you specify that the range is to be all numbers between 0 and 5. Call the random number generated Y. Then the density curve of the random variable Y has constant height between 0 and 5, and height 0 elsewhere.

(a) What is the height of the density curve between 0 and 5? (Enter your answer to two decimal places.)

c) Find

P(0.5 < Y < 1.8).

(Enter your answer to three decimal places.)
Answer #1

a) Height of the density curve = 1/(5 - 0) = 1/5 = 0.20

c) P(0.5 < Y < 1.8) = (1.8 - 0.5) / (5 - 0) = 1.3/5 = 0.26